The total number of hours, measured in units of 100 hours, that a family runs a vacuum cleaner over a period of one year is a continuous random variable X that has the density function.


f(x) = x, 0 < x < 1,

2 − x, 1 ≤ x < 2,

0, elsewhere.


Find the probability that over a period of one year, a

family runs their vacuum cleaner

(a) less than 120 hours;

(b) between 50 and 100 hours.


1
Expert's answer
2022-02-11T03:18:28-0500

a)


"x = 120\/100 = 1.2"


"P (x < 1.2) = \int^1_0xdx+\int^{1.2}_1(2-x)dx"

"=[\dfrac{x^2}{2}]\begin{matrix}
 1 \\
 0
\end{matrix}+[2x-\dfrac{x^2}{2}]\begin{matrix}
 1.2 \\
 1
\end{matrix}"

"=0.5+2.4-0.72-2+0.5=0.68"

b)


"x = 50\/100 = 0.5,"

"x = 100\/100 = 1"

"P (0.5<x < 1) = \int^{1}_{0.5}xdx=[\dfrac{x^2}{2}]\begin{matrix}
 1 \\
 0.5
\end{matrix}"

"=0.5-0.125=0.375"